Calculate the future value of end-of-month payments of $3,000 made for 6 years into an investment fund that earns 6.50% compounded quarterly.
Answer #1
Quarterly rate = 6.5%/4 =1.625%
Let us assume monthly effective rate r
equation would be
1.625% =(1+r)^3 -1
r =0.5387918%
Future Value of an Ordinary Annuity
= C*[(1+i)^n-1]/i
Where,
C= Cash Flow per period
i = interest rate per period =0.5387918%
n=number of period =6*12 =72
= $3000[ (1+0.005387918)^72 -1] /0.005387918
= $3000[ (1.005387918)^72 -1] /0.005387918
= $3000[ (1.4724 -1] /0.005387918]
= $2,63,028.89